If demand function is given as the following: 

Qz = 230 -2.75 Pz + 0.5 I + 1.2 Pm + 0.6A 

Where Qz is quantity of Good z sold, Pz is price of Good z per unit, I is per capita income, Pm is price of competitor and A is the amount of advertising spent. 

Current values:  Pz= RM 55 I= RM 9000    Pm= RM 50    A =RM 12,000

a)   Should the firm consider giving a price discount in order to increase total revenue?



1
Expert's answer
2021-05-03T10:55:10-0400
"Q_Z=230-2.75\times55+0.5\times9000+1.2\times50+0.6\times12000"

"Q_Z=11990"

"E=-2.75\times\frac{55}{11990}=-0.01"

As you can see, in this case, absolutely inelastic demand, that is, demand, the volume of which remains unchanged for any price changes. Consequently, the firm can increase the price of the product to generate more revenue.